# Generalized Bayesian deep reinforcement learning

### v1

@@ A.2 Approximation error of the Q function
$\leq\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}K_{\mu_{j+1}}(s,a)$
@@
$=\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\gamma||Q^{\bm{\theta^{(k)}}}_{\mu_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{\gamma}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j}}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|(\gamma K^{\prime}_{\mu_{j}}+K^{\prime}_{\mu_{j+1}})$
@@
where $\zeta_{j}(k,n)=||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{\gamma}||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}|\theta_{ki}-\theta^{*}|(K^{\prime}_{\mu_{j}}+\frac{1}{\gamma}K^{\prime}_{\mu_{j+1}})$
Let us define, $Y_{k}=\theta-\hat{\theta_{k}}$, where $\theta\sim\pi_{k}$ and $\hat{\theta_{k}}$ denotes the Scoring rule minimizer obtained after observing $k$th episode. So, according to the Bernstein–von Mises (BvM) theorem, $Y_{k}=O_{p}((k\tau)^{-1/2})$ assuming $Y_{k}$ has finite expectation. Furthermore, from the consistency results of M-estimators (Van der Vaart, 2000), under certain regularity conditions,$\hat{\theta_{k}}$ converges to the expected Scoring Rule minimizer ($\theta^{*}$) at the rate of $O_{p}((k\tau)^{-1/2})$.
For the $n$ samples drawn from the $k$th posterior, if we define $Y_{ki}=|\theta_{ki}-\hat{\theta_{k}}|+|\hat{\theta_{k}}-{\theta^{*}}|$ then, $Y_{ki}=O_{p}((k\tau)^{-1/2})$ for all $i=1,2,\ldots n$. Hence, $1/n\sum_{i=1}^{n}|Y_{ki}|=O_{p}((nk\tau)^{-1/2})$. Therefore, the last term in the equation (A.2) vanishes with large $k$ and $n$. Thus, compared to the classical TS, for ETS the last term shrinks $\sqrt{n}$ times faster.
Note that, $\mu^{*}_{j}(s)=\arg\max_{a}Q_{\mu^{*}_{j-1}}^{\theta^{*}}(s,a)$ and $\mu_{j}(s)=\arg\max_{a}Q_{\mu_{j-1}}^{\bm{\theta(k)}}(s,a)$. So for large enough $k$ and $n$, from the consistency of posterior mean, we can say that, $Q_{\mu^{*}_{j}}^{\theta^{*}}\approx Q_{\mu_{j}}^{\bm{\theta(k)}}$ for any policy iteration step $j$. Then, the second and third term in equation (A.2) also vanish as $k\rightarrow\infty$ and $n\rightarrow\infty$.
A.3 Tuning parameters for posterior sampling using SMC
@@
A.4 Zeroth order gradient
@@
A.5 Implementation details of REINFORCE

### v2

@@ Appendix B Approximation error of the Q function
$\leq\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}K_{\mu_{j+1}}(s,a)$
@@
$=\gamma||Q^{*}-Q^{\theta^{*}}_{\mu^{*}_{j}}||_{\infty}+||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\gamma||Q^{\bm{\theta^{(k)}}}_{\mu_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{\gamma}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||K^{\prime}_{\mu_{j}}+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||K^{\prime}_{\mu_{j+1}}$
@@
$\quad+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||(\gamma K^{\prime}_{\mu_{j}}+K^{\prime}_{\mu_{j+1}})$
@@
where $\zeta_{j}(k,n)=||Q^{\theta^{*}}_{\mu^{*}_{j}}-Q^{\theta^{*}}_{\mu_{j}}||_{\infty}+\frac{1}{\gamma}||Q^{\theta^{*}}_{\mu^{*}_{j+1}}-Q^{\theta^{*}}_{\mu_{j+1}}||_{\infty}+\frac{1}{n}\sum_{i=1}^{n}||\theta_{ki}-\theta^{*}||(K^{\prime}_{\mu_{j}}+\frac{1}{\gamma}K^{\prime}_{\mu_{j+1}})$
Let us define, $Y_{k}=\theta-\hat{\theta}_{k}$, where $\theta\sim\pi_{k}$ and $\hat{\theta}_{k}$ denotes the Scoring rule minimizer obtained after observing $k$th episode. So, according to the Bernstein–von Mises (BvM) theorem, $Y_{k}=O_{p}((k\tau)^{-1/2})$ assuming $Y_{k}$ has a finite expectation. Furthermore, from the consistency results of M-estimators (Van der Vaart, 2000), under certain regularity conditions,$\hat{\theta}_{k}$ converges to the expected scoring rule minimizer ($\theta^{*}$) at the rate of $O_{p}((k\tau)^{-1/2})$.
For the $n$ samples drawn from the $k$th posterior, if we define $Y_{ki}=||\theta_{ki}-\hat{\theta_{k}}||+||\hat{\theta_{k}}-{\theta^{*}}||$ then, $Y_{ki}=O_{p}((k\tau)^{-1/2})$ for all $i=1,2,\ldots n$. Hence, $1/n\sum_{i=1}^{n}|Y_{ki}|=O_{p}((nk\tau)^{-1/2})$. Therefore, the last term in the equation (B) vanishes with large $k$ and $n$. Thus, compared to the classical TS, for ETS, the last term shrinks $\sqrt{n}$ times faster.
Note that, $\mu^{*}_{j}(s)=\arg\max_{a}Q_{\mu^{*}_{j-1}}^{\theta^{*}}(s,a)$ and $\mu_{j}(s)=\arg\max_{a}Q_{\mu_{j-1}}^{\bm{\theta^{(k)}}}(s,a)$. So for large enough $k$ and $n$, from the consistency of posterior mean, we can say that, $Q_{\mu^{*}_{j}}^{\theta^{*}}\approx Q_{\mu_{j}}^{\bm{\theta^{(k)}}}$ for any policy iteration step $j$. Then, the second and third term in equation (B) also vanish as $k\rightarrow\infty$ and $n\rightarrow\infty$.
Appendix C Implementation details
C.1 Tuning parameters for posterior sampling using SMC
We provide a list of values of the tuning parameters used for the SMC sampler in Table 1.
@@ C.1 Tuning parameters for posterior sampling using SMC
C.2 Zeroth order gradient
@@
C.3 Implementation details of REINFORCE